OpenStudy (anonymous):

x^2-10x=-21 solve by factoring

OpenStudy (anonymous):

I want to say 3,7

OpenStudy (haseeb96):

x^2-10x+21=0 x^2-3x-7x+21=0 x(x-3)-7(x-3)=0 (x-7)(x-3)=0 either x-7=0 or x-3=0 x=7 or x=3 this is the answer